# Seewo’s Submission to MLC-SLM: Lessons learned from Speech Reasoning Language Models

### v1

@@ Abstract
This paper presents Seewo’s systems for both tracks of the Multilingual Conversational Speech Language Model Challenge (MLC-SLM), addressing automatic speech recognition (ASR) and speaker diarization with ASR (SD-ASR). We introduce a multi-stage training pipeline that explicitly enhances reasoning and self-correction in speech language models for ASR. Our approach combines curriculum learning for progressive capability acquisition, Chain-of-Thought data augmentation to foster intermediate reflection, and Reinforcement Learning with Verifiable Rewards (RLVR) to further refine self-correction through reward-driven optimization. This approach achieves substantial improvements over the official challenge baselines. On the evaluation set, our best system attains a WER/CER of 11.57% for Track 1 and a tcpWER/tcpCER of 17.67% for Track 2. Comprehensive ablation studies demonstrate the effectiveness of each component under challenge constraints.
Index Terms: multilingual speech recognition, speech language model, speech diarization, self-correction, reinforcement learning.
@@ 3 Training pipeline of ASR system
We employ a multi-stage training approach, where each stage addresses specific challenges encountered during model development. The latter stage is continued fine-tuning on the checkpoint of the previous one.
@@ 3.3 Stage 3: LoRA training with multilingual prompts and context
To address these issues, we adopted LoRA [23] adaptation on both the LLM decoder and Whisper encoder. We configured the LoRA modules with rank 16 and lora_alpha 32, providing a balance between model capacity and training efficiency. Additionally, we introduced language-specific prompts in the <P> section, which were ”Transcribe the speech above” translated into each respective language.
@@ 3.4 Stage 4: SFT for reflection
We hypothesize that the standard loss function is suboptimal for this task, primarily due to two factors: (1) the completion length $|I_{c}|$ inversely affects the loss magnitude, and (2) the <think> section, which in most cases is twice as long as the <transcribe> section, disproportionately influenced the loss. Recent studies [26, 7] have shown that the completion-to-prompt length ratio ($R_{g}$) significantly impacts model performance. In our experiments, this imbalance resulted in close loss values for stages 3 and 4, despite substantial differences in WER.
@@ 3.5 Stage 5: RLVR for reflection
For the $RF_{3}$, we observed that the model’s output was getting more accurate with the error type identification. The reward of this function converged above 0.6; however, while this did not translate into improved WER/CER. The process reward did not lead to WER/CER improvement in our experiments. For the $RF_{4}$, we found that the model was not able to generate the correct error details, the reward of this function converged to 0.5. We conjectured that the reward signal was too sparse to guide the model to generate the correct error details. Until the end of the challenge, we did not find a good way to improve the self-correction ability of the model, which would be a future work.
Besides, we found that it was important to obtain the capability step by step with a curriculum learning strategy. In the attempt of $RF_{2}$+$RF_{5}$, the <think> structure format was collapsed quickly, and the rewards of $RF_{2}$ and $RF_{5}$ kept in a low level. The GRPO reinforcement learning process was trying to getting the better sampling signal, the premise is that better sampling results exists in a group. [9]
@@ 4 Additional Experiments
During the evaluation phase of the challenge, we also found some other approaches to improve the performance of the ASR task.
@@ 4.1 Effect of Conversational Context on ASR
Inspired by GEC-RAG [27], which demonstrated that contextual information can enhance error correction in post-processing, we explored the impact of conversational context on ASR model performance in the MLC-SLM challenge. Given the conversational and long-form nature of the audio, we hypothesized that incorporating preceding utterances could improve the model’s self-correction ability. During training, we augmented each sample by appending the transcriptions of the previous two utterances in the same conversation to the user prompt. With the new prompt, we fine-tuned the model from ckpt-7000 of Stage 3 to ckpt-9000. At inference time, the model was also provided with the transcriptions of the previous two utterances as context. The results are summarized in Table 3.
@@
ckpt -9000 (with context) 0.14 14.30%
It should be noted that, in this experiment, the contextual transcriptions used during evaluation were taken from the ground truth of the development set, rather than generated by the model itself. As a result, the reported results represent an upper bound on the potential benefit of context, rather than the true end-to-end performance. Due to the inefficiency of our inference pipeline, which is slow for sequential utterance processing by the conversational order, we were unable to conduct large-scale, fully automatic experiments, fully automatic experiments at scale. Improving the inference efficiency and conducting stricter end-to-end evaluations will be the focus of our future work.
@@ 6 Conclusions
In this work, we present a comprehensive multi-stage training pipeline for multilingual conversational ASR and speaker diarization, integrating curriculum learning, Chain-of-Thought data augmentation, and RLVR-based self-correction. Our approach achieves substantial improvements over the official challenge baselines. We summarize the results of our main version submitted to the evaluation set in table 6 and 7, to demonstrate the effectiveness of each component. Despite these advances, challenges remain in fully automating context-aware inference and further improving the granularity of self-correction. Future work will focus on enhancing inference efficiency, developing more sophisticated reward functions for reasoning, and extending our methods to broader real-world multilingual and conversational scenarios.
Table 6: Results submitted to the evaluation set of track 1
@@
ckpt-12500 RLVF($RF_{1}$+$RF_{2}$) 13.15% 11.79%
ckpt-12500 RLVF($RF_{1}$+$RF_{2}$+$RF_{5}$) 12.73% 11.57%
Table 7: Results submitted to the evaluation set of track 2

### v2

@@ Abstract
This paper presents Seewo’s systems for both tracks of the Multilingual Conversational Speech Language Model Challenge (MLC-SLM), addressing automatic speech recognition (ASR) and speaker diarization with ASR (SD-ASR). We introduce a multi-stage training pipeline that explicitly enhances reasoning and self-correction in speech language models for ASR. Our approach combines curriculum learning for progressive capability acquisition, Chain-of-Thought data augmentation to foster intermediate reflection, and Reinforcement Learning with Verifiable Rewards (RLVR) to further refine self-correction through reward-driven optimization. Our approach substantially outperforms the official baselines, achieving 11.57% WER/CER (Track 1) and 17.67% tcpWER/tcpCER (Track 2) on the evaluation set. Comprehensive ablation studies validate each component’s effectiveness.
Index Terms: multilingual speech recognition, self-correction, reinforcement learning, speech language model, speech diarization
@@ 3 Training pipeline of ASR system
We adopt a multi-stage training pipeline where each stage builds upon the previous checkpoint to address specific challenges. While we present only the successful experiments here, our development process included numerous exploratory attempts that, despite not improving performance, provided valuable insights into the model’s capabilities and limitations.
@@ 3.3 Stage 3: LoRA training with multilingual prompts and context
To address these issues, we adopted LoRA [23] adaptation (with rank 16 and lora_alpha 32) on both the LLM decoder and Whisper encoder. Additionally, we introduced language-specific prompts in the <P> section, which were ”Transcribe the speech above” translated into each respective language.
@@ 3.4 Stage 4: SFT for reflection
We hypothesize that the standard loss is suboptimal due to (1) completion length $|I_{c}|$ inversely affecting loss magnitude, and (2) the typically longer <think> section disproportionately influencing the loss. Recent studies [26, 7] have shown that the completion-to-prompt length ratio ($R_{g}$) significantly impacts model performance. In our experiments, this imbalance resulted in close loss values for stages 3 and 4, despite substantial differences in WER.
@@ 3.5 Stage 5: RLVR for reflection
For the $RF_{3}$, we observed that the error type identification was getting more accurate during training, and the reward of $RF_{3}$ converged above 0.6; however, this did not translate into improved WER/CER. For the $RF_{4}$, we found that the model was not able to generate the correct error details, the reward of this function converged around 0.5. We conjectured that the reward signal was too sparse to guide the model to generate the correct error details.
In our experiments, the outcome-based reward functions were found to be an effective strategy for training models to perform self-correction, even though the intermediate reasoning steps did not demonstrate a direct causal influence on the final output. This observation aligns with recent findings in the literature [27] [28] that suggest the process-based reward functions might be less stability and need more refinement. By the end of the challenge we had not found a good way to improve the reflection ability of the model, which would be a future work.
Furthermore, we found that it was important to obtain the capability step by step with a curriculum learning strategy, the model’s low level capability must be established first. In the attempt of $RF_{2}$+$RF_{5}$, the <think> structure format was collapsed quickly, and the rewards of $RF_{2}$ and $RF_{5}$ remained low. The GRPO reinforcement learning process was trying to get the better sampling signal, the premise is that better sampling results exist in a group. [9]
@@ 4.1 Effect of Conversational Context on ASR
Inspired by GEC-RAG [29], which demonstrated that contextual information can enhance error correction in post-processing, we explored the impact of conversational context on ASR model performance in the MLC-SLM challenge. Given the conversational and long-form nature of the audio, we hypothesized that incorporating preceding utterances could improve the model’s self-correction ability. During training, we augmented each sample by appending the transcriptions of the previous two utterances in the same conversation to the user prompt. With the new prompt, we fine-tuned the model from ckpt-7000 of Stage 3 to ckpt-9000. At the test time, the previous two utterances were used as context as well. The results are summarized in Table 3.
@@
ckpt -9000-context (with context) 0.14 14.30%
It should be noted that, in this experiment, the contextual transcriptions used during evaluation were taken from the ground truth of the development set, rather than generated by the model itself. As a result, the reported results represent an upper bound on the potential benefit of context, rather than the true end-to-end performance. Due to the inefficiency of our inference pipeline, which is slow for sequential utterance processing by the conversational order, we were unable to conduct large-scale, fully automatic experiments at scale. Improving the inference efficiency and conducting stricter end-to-end evaluations will be the focus of our future work.
@@ 6 Conclusions
In this work, we present a comprehensive multi-stage training pipeline for multilingual conversational ASR, integrating LoRA, Chain-of-Thought data augmentation, and RLVR-based self-correction. Our approach achieves substantial improvements over the official challenge baselines. We summarize the results of our main version submitted to the evaluation set in table 6 and 7, to demonstrate the effectiveness of each component. Despite these advances, challenges remain in fully automating context-aware inference and further improving the granularity of self-correction. Future work will focus on enhancing inference efficiency, developing more sophisticated reward functions for reasoning, and extending our methods to broader real-world multilingual and conversational scenarios.
Table 6: Results of development and evaluation sets of track 1
@@
ckpt-12500-A RLVF($RF_{1}$+$RF_{2}$) 13.15% 11.79%
ckpt-12500-B RLVF($RF_{1}$+$RF_{2}$+$RF_{5}$) 12.73% 11.57%
Table 7: Results of development and evaluation sets of track 2